Select any two integers between -12 and +12 which will become solutions to a system of two equations.
- The two integers I picked are -4 and 8
Write two equations that have your two integers as solutions. Show how you built the equations using your integers.
- The two equations I picked are 5X-Y=-4 and 6X+4Y=8.
Solve your system of equations by the addition/subtraction method. Make sure you show the necessary 5 steps.
- Both equations in ax + by= c is written as 5x-y=-4 and 6x+4y=8.
Step 1: Write both equations in the form ax + by = c
Step 2: Multiply the second equation by 2 in order to make the coefficients of the x terms equal.
Step 3: Subtract the second equation from the first equation to eliminate the x variable.
Solve the equation for y.
Step 5: Select one equation and substitute 4 for y and solve for x.
The solution set are?
Please help with this.

Answer by richwmiller(17219)

The solution to the two equations that you chose are not -4 and 8 but X = -4/13, Y = 32/13
why do you think that -4 and 8 would be the solutions?
5X-Y=-4
5*4-8=a
-20-8=-28
5X-Y=-28
6X+4Y=8
and 6(-4)+4*8=b.
-24+32=8
6X+4Y=8
which is the original equation you had
6X+4Y=8,
5X-Y=-28
now the solution is x=-4 and y=8
(-4, 8)